OpenStudy (anonymous):

integrate(-2cos(x)/[y*sin(x)])dx

ganeshie8 (ganeshie8):

is that part of a double integral ? whats `y` doing in the denominator ?

OpenStudy (anonymous):

it's not a double integral, this integral came out as a intermediate step in solving a differential integral

OpenStudy (anonymous):

I'm not sure how to treat the y in the denominator. I was thinking of treating it as a constant?

OpenStudy (anonymous):

solving a differential equation*, not differential integral* typo

ganeshie8 (ganeshie8):

thats not allowed then whats the original differential equation ?

OpenStudy (dumbcow):

yeah you have to completely separate variables before integrating

OpenStudy (anonymous):

\[\frac{ 3y ^{2}\cot(x)+\sin(x)\cos(x) }{ 2y } = \frac{ dy }{ dx }\]

OpenStudy (dumbcow):

looks like a bernoulli equation http://tutorial.math.lamar.edu/Classes/DE/Bernoulli.aspx

OpenStudy (anonymous):

Oh it does look like a bernoulli equation. We haven't covered it in class yet but I have seen it on the course outline.

OpenStudy (dumbcow):

oh ok, well there may be another way , what are you covering right now?

OpenStudy (anonymous):

1. Direct Integration 2. Substitution --> Integration 3. Exact Equations 4. Made Exact - Integrating Factor 5. First Order Linear Fomula 6. Linearlize --> Use formula

OpenStudy (dumbcow):

you are going to make a substitution to change it to a Linear DE Bernoulli is just a nice general form to help tell you exactly what substitution to make not sure what your teacher wants exactly, prob covered in topic 5 and 6 i guess

OpenStudy (dumbcow):

substitution is: v = y^2 v' = 2y y' just say you skipped to bernoulli section haha

OpenStudy (anonymous):

haha thanks, I'll try it

OpenStudy (anonymous):

Just a quick question though about the two variable integral i mentioned earlier;

OpenStudy (anonymous):

so you can't integrate when there is x and y in the function if it's just integrating by dx, but I know from other calc courses that if it's a double integral,

OpenStudy (anonymous):

you can integrate with respect to dxdy and treat one of the variables constant as you integrate.

OpenStudy (anonymous):

So you can't apply this procedure of keeping variables constant in the integral i mentioned above?

OpenStudy (anonymous):

sorry about reply being broken up^ the post button didnt' work when it was all together ..

OpenStudy (dumbcow):

hmm well to use a double integral you need dxdy to be a product on same side that usually doesn't work out in most DE

OpenStudy (dumbcow):

how do you turn dy/dx into dy*dx ?

OpenStudy (dumbcow):

also with a double integral you need limits

OpenStudy (gorv):

can u send the Q which is given and you have to solve

OpenStudy (anonymous):

The original ODE problem is: \[\frac{ 3y ^{2}\cot(x)+\sin(x)\cos(x) }{ 2y }\] But I solved it, thanks